OpenStudy (anonymous):

What is the equation of a parabola with a vertex at (-2, 0) and a directrix x = -4.

OpenStudy (ranga):

The distance between the vertex and the directrix is same as the distance from the vertex to the focus. Distance from vertex to directrix = -2 - (-4) = -2 + 4 = 2 To find the focus, add 2 to the vertex: Focus is at (-2+2, 0) or (0,0) A parabola has the property that all points on the parabola are equal distance from the focus and the directrix. Let (x,y) be a general point on the parabola. The square of its distance from the focus is: (x - 0)^2 + (y - 0)^2 = x^2 + y^2 --- (1) The distance of (x,y) from the directrix x = -4 is: (x - (-4)) = x + 4. Square this distance and equate it to (1) x^2 + y^2 = (x + 4)^2 expand the right and simplify.

OpenStudy (ranga):

x^2 + y^2 = (x + 4)^2 x^2 + y^2 = x^2 + 8x + 16 y^2 = 8x + 16 y^2 - 16 = 8x x = 1/8y^2 - 2 This is a horizontal parabola that opens to the right and its equation is x = 1/8y^2 - 2

OpenStudy (anonymous):

Thank you so much you helped a lot!

OpenStudy (ranga):

You are welcome.